OpenStudy (anonymous):

Solve the following system of equations. 2x – y + z = –3 2x + 2y + 3z = 2 3x – 3y – z = –4 Answer (1, 3, –2) (1, 3, 2) (1, –3, 2) (–1, 3, 2)

OpenStudy (anonymous):

2x-y+z=-3 3x-3y-z=-4 => 5x-4y=-7 (1) 2x+2y+3z=2 3x-3y-z=-4 =>11x-7y=-10 (2) (1) and (2) => 5x-4y=-7 11x-7y=-10 => 9x=9 => x=1 y=3 z= -2 (1, 3, -2)